UT 20120203 AGw#4 at RFBG Verification

Observer: DMiller, JCarlos, JHill (LBTO)
Telescope Operator: DHuerta (LBTO)
Telescope Support: (LBTO)
Instrument Support: DOfficer (LBTO)

Summary

Details

4:30 Open Chamber

4:35 John tried to set the left Adsec, but Arbitrator was down. Was down since Italians were testing with AGW#1. Juan Carlos started AO arbitrator.

4:38 John set left Adsec. Success

4:49 preset BS9122 mode=ACTIVE, GS=0. Failed. W-Unit at LBBG software is running, but with some systems off. Alfio Fix by restarting the W-Arbitrator. Before W-Arbitrator thought the W-Unit was on, but was off, and could not stop the W-Unit. Restarting the W-Arbitrator made it realize the W-unit was powered off.

4:51 preset again and failed because could not move guide probe to position. Looks like AGw is failing to home. John tries to stop and start AGw#3.

5:05 reboot agw0 computer. Still the same problem: AGw#3 motor 1 end limit set, crash limit encountered....

Andrew's Laser Tracker Measurements

5:30 Go to Zenith and reconfigure to MODS to collect Andrew's laser tracker data.

5:40 have to turn on MODS agw and open hatch

6:00 Could not find M5 star. Spiral search. Still no star. Rehome MODS agw. Star is not where it should be.

6:15 preset BS9122 MODE=ACTIVE GS=0

6:30 preset BS9126

6:45 preset BS9148 and BS9145. Andrew's laser tracker was pointing near the secondary and causing glints.

7:00 preset BS91...5 more points

Coma test for guiding jumps

8:06 preset BS9142: 3'+ seeing. Can not collimate. Seeing is too bad for this test even!

Pisces filter wheel

8:19 JC is going to power off the Pisces controller, I will reboot the Pisces Computer, JC will turn the power on the controller and press the reset button. I will then startup and activate Pisces using my IDL. The hope is to get the filter wheel to move. It did not early when JC tried.

9:00 Success. Full power off of computer and power off of filter wheel controller has fixed the problem. JC found a problem (maybe) with the filter wheel positions. At least they are not what we expected.

10:00 Preset BS9142 at 77 elevation. Seeing about 1.3"

10:15 preset BS9162 at 40 elevation for Andrew's data set. Seeing 2.0"

10:25 preset BS9169 at 30 elevation for Andrew's data set. Seeing 2.1" FWHM but should be well enough collimated for Andrew.

Coma test

Coma Z7 10:39:28, 40:58, 41:33, 42:25 +-500, +-1000

Coma Z8 10:43:59, 45:14, 45:35 +-500 +-1000

Focus z4 10:47:21, 47:48, 48:11 +-1000

Could not see any big jumps during the application of these aberrations. Turned off guiding and collimation for all sets.

Alignment of ND holder

11:00 JC take blank and J images. Can see 90 degree corner in upper left quadrant. pisces image 39, 40

11:20 JC remove holder. Took more image with and without holder and in two filter, but not sure we were seeing what we needed.

-- DougMiller - 31 Jan 2012
Topic revision: r5 - 03 Feb 2012, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback